OpenStudy (brooke2353):

A physician saw 18 cases of chicken pox in the first week of an outbreak. Each week, he saw 6 more cases than the week before. After 10 weeks, how many cases of chicken pox did the physician see altogether?? A. Write a sequence that shows the number of cases seen for each of the 10 weeks during the outbreak. B. Use the Gauss' method to find the sum. Be sure to show and explain your work. plz help me

OpenStudy (anonymous):

18...24...30...and so on and so forth. All you must do is add six to the original number for each week. I can't remember what Guass' formula is though.

OpenStudy (anonymous):

\[\sum_{k=1}^{10}(12+6k)=\sum_{k=1}^{10}12+6\sum_{k=1}^{10}k=120+330=450\]Gauss' method solves the sum of the first n numbers. The formula used here is \[\sum_{k=1}^{n}k=\frac{n(n+1)}{2}\]